A mixture of birdseed contains sunflower seeds worth $1.30 per kilogram and other seeds worth #0.90 per kilogram. How many kilograms of each kind must be used to produce 100 kilograms of mixture that sells $1.00 per kilogram?
Answer by ptaylor(2198)

Let x=amount of $0.90 seeds needed; $0.90x=value of these seeds
Then 100-x=amount of $1.30 seeds needed; $1.30(100-x)= value of these seeds
And $1.00(100) equals value of final mixture
Now we know that the value of the seeds before the mixture takes place has to equal the value of the seeds after they are mixed together. So, our equation to solve is:
$0.90x+$1.30(100-x)=$1.00*100 get rid of parens (distributive law)
$0.90x+$130-$1.30x=$100 subtract $130 from each side
$0.90x+$130-$130-$1.30x=$100-$130 collect like terms
-$0.40x=-$30 divide each side by -$0.40
x=75 kilo-------------------------------------amount of $0.90 seeds needed
100-x=100-75=25 kilo ------------------------amount of $1.30 seeds needed
CK
75*$0.90+25*$1.30=100*$1.00
$67.5+$32.5=$100
$100=$100
Hope this helps---ptaylor